DX Reconstructor Measurements

17:48
  • Waiting for the telescope to come up on oil and the lights to be turned off.
  • Applied the flat 20190208_400m_600hz_1.sav
  • 634 modes.
18:20
  • Found the light, everything running.
  • When the latest recon was applied the mirror tipped until the light was lost in the pupils
  • Could be due to the Tip-tilt mirror not functioning
  • Tip-tilt mirror is not working, stand-by, re-starting WFS
  • Saved Board Setup: 2019_0208
18:36
  • Restarted W-unit, it seems to be working OK, now
  • Pupils look nice and round
  • 600Hz 3 lambda/d
  • Disturbance, using: doug_400_2l_0.6.fits
  • Disturb: 20190208_184159
  • Away we go!
  • 4 iteration Intmat, closed loop with Rec_20190208_112649.fits
  • Intmat: 20190208_184405
  • Recon: 20190208_184602
  • Looks OK, Nice and stable, but I still think that the TT mirror isn't working right, the pupils flash no matter how much TT gain I apply
  • That's not quite true, the TT has a little effect, I think that taking another 4 iteration reconstructor at 3 lambda/d will help, let's see
  • Intmat: 20190208_185639
  • Recon: 20190208_185800
19:06
  • Made the AdSec go into failure because I increased the TT gain to 0.65, without any high order gains, this makes the adsec crash
  • TT correction seems to be working, optimizing gain.
  • 1.23 0.1 0.1 optimized gain
  • Going to take a 12 iteration recon, in order to get a better recon, before moving on to a long series.
  • Intmat: 20190208_191430
  • Recon: 20190208_191824
  • Small bite at 7:00 position, but much better than I've seen, overall.
19:22
  • Let's beat this sucker to a pulp
  • 96 Iterations
  • Intmat: 20190208_192454
  • Recon: 20190208_195120
  • Something Odd happened and the pupils seem to have an odd variation, in a few modes.
  • The gain plot, during the analysis showed an odd and very large peak around mode 200
  • 48 Mode recon, to see if things improve
  • Intmat: 20190208_200132
  • Recon: 20190208_201527
20:17
  • Time to test!
  • Test W/O Disturbance looks good
  • Disturb: /local/aomeas/adsec_calib/CMD/disturb/atm_0.8seeing_1000Hz.fits
  • Looks beautiful
  • Modal Plot: 202415
  • Oops, we're still at 600Hz
  • 990Hz
  • Wow, that's smooth!
  • Modal Plot: 202656
  • The modal plot is great!
  • Both at 600 and 990Hz
  • Moving along, let's try atm_1.6seeing_1000Hz.fits
  • I realize that I left the camera lens off, after optimizing the gain, let's see if that makes a difference
  • 203108
  • Fan-bleeding-tastic.
20:32
  • I have to do the other binnings, there isn't much time.
  • Bin4 might let me start from the good flat and the previous bin4 reconstructor
  • Lamp intensity 32
  • No, there are some weird features in the pupil, it's not as smooth as it should be.
  • I have to take out the coma.
  • I can probably start with the 10 mode bin 4 recon, that will take out the coma
  • Going back to bin 1 to flatten it, stupid thing.
  • Done
  • Bootstrapping from open loop
  • Disturb: 20190208_205518
  • Intmat: 20190208_205605
  • Recon: Failed! (Did only 1 iteration, this is what happens when you are in a hurry)
  • Intmat: 20190208_205933
  • Recon: 20190208_210057
  • Looks OK, now in closed loop
  • Shell went "not Flat" when I opened the loop (scrunched the pupils)
  • Nothing else for it, going back to bin 1 to re-flatten (didn't save a bin1 flat, because I'm dumb)
  • OK, much better.
  • Intmat: 20190208_210937
  • Recon: Failed!? This time I did 4 iterations, oh well (at 600Hz with the correct pupils)
  • Intmat: 20190208_211228
  • Recon: 20190208_211352
21:14
  • 36 mode, 4 iters
  • Disturb: 20190208_211517
  • Intmat: 20190208_211650/
  • Recon: 20190208_211821
  • Looks OK
  • After centering the pupils, I'll take a long series
21:19
  • 48 Iterations
  • Intmat: 20190208_212057
  • Recon: 20190208_213433
  • Fan-bleeding-tastic
  • Optical Loop Whatsis: 213944
  • Hee Hee, no disturbance
  • Going straight to the 1.6" disturb.
  • Can't do TT modulation of 6 at 990Hz, sticking to 600
  • 214430
  • Good correction up to mode 36
21:45
  • To Bin2, if we still have time
  • New Bin1 flat: 20190208_400m_600hz_2
  • Now to bin2
  • Boot-strap from 10 modes
  • Disturb: 20190208_215203
  • Intmat: 20190208_215450
  • REcon: 20190208_215609
  • Stable, centering pupils
  • 10 Mode Closed Loop
21:58
  • Shell ripped when resetting the flat, jumping actuator pushed forces way out of bounds?
  • D*** it! Don't have time for this
  • Panic, doing quick restart (IDL, fastdiagn, arbitrator)
22:05
  • It set with 634 modes, I don't know what happened, the Adsec must have been grumpy
  • Intmat: 20190208_220722
  • Recon: 20190208_220852
  • Looks good to me
22:09
  • 36 Modes
  • Disturb: 20190208_221058
  • Intmat: 20190208_221137
  • Recon: 20190208_221341
  • Looks fine
22:14
  • 66 Modes
  • Disturb: 20190208_221548
  • Intmat: 20190208_221625
  • Recon: 20190208_221805
  • Stable
22:21
  • 153 Modes
  • Disturb: 20190208_222059
  • Intmat: 20190208_222224
  • Recon: 20190208_222348
  • This one is behaving weirdly, significant oscillation in the mid range modes, it looks like
  • Trying again
  • Closed loop with 66 modes
  • Intmat: 20190208_222614
  • Recon: Failed! (Why do these things keep failing)
  • Intmat: 20190208_223033
  • Recon: 20190208_223254
  • Much Better
22:34
  • 48 Iterations
  • Intmat: 20190208_223619
  • Recon: 20190208_224914
  • GRR! It's oscillating again, al;skdjfa;lskdfj;alskjdf;la!!
  • I did 49 iterations, does that have something to do with it?
  • Intmat: 20190208_225609
  • Recon: 20190208_230910
  • Good, no funny oscillation
  • Test with 1.6" seeing
  • Looks very stable
  • Optical loop data, 600 Hz: 231410
  • Looks OK
  • Optical loop data, 990 Hz: 231638
  • The last 25 modes don't look that hot, they do correct, but not as well as exped, but I'm out of time for bin2
23:18
  • Bin 3 ahoy!
  • 10 modes open loop
  • Disturb: 20190208_232203
  • Intmat: 20190208_232230
  • Recon: 20190208_232350
  • Closed loop 10 modes
  • Intmat: 20190208_232550
  • Recon: 20190208_232713
  • Stable
23:30
  • 36 Modes
  • Disturb: 20190208_232952
  • Intmat: 20190208_233111
  • Recon: 20190208_233230
  • Nice and stable
23:33
  • 66 Modes
  • Disturb: 20190208_233444
  • Intmat: 20190208_233529
  • Recon: 20190208_233647
  • Looks Pretty Good
23:38
  • 48 iterations
  • Intmat: 20190208_234447
  • Recon: 20190208_235749
  • Test with 1.6" seeing, I'm outta time
  • 000138
  • Looks pretty good!
  • Finished
  • I'll calculate the bad seeing and 12 mode recons.
  • 400 mode, bin1: 20190208_201527
  • 300 mode, bin1: 20190209_002624
  • 12 mode, bin1: 20190209_002846
  • 153 mode, bin2: 20190208_230910
  • 110 mode, bin2: 20190209_003800
  • 12 mode, bin2: 20190209_004016
  • 66 mode, bin3: 20190208_235749
  • 12 mode, bin3: 20190209_004158
  • 36 mode, bin4: 20190208_213433
  • 12 mode, bin4: 20190209_003241

-- GregoryTaylor - 08 Feb 2019
Topic revision: r1 - 09 Feb 2019, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback